OpenStudy (leenathan):

help plz will fan and medal What is the volume of a cone with a radius of 3 feet and a height of 6 feet? Use 3.14 for pi. Round your answer to the nearest hundredth. 18.31 ft3 40.25 ft3 56.52 ft3 60.34 ft3

OpenStudy (studygurl14):

Volume of cone \(\Large\frac{1}{3}\pi r^2 h\)

OpenStudy (leenathan):

a?

OpenStudy (studygurl14):

no. plug in known values ino the formula

OpenStudy (leenathan):

oh ok

OpenStudy (leenathan):

ummm how?

OpenStudy (studygurl14):

r = 3 h = 6 pi = 3.14

OpenStudy (leenathan):

ok so 13*3.14*3*2*6

OpenStudy (leenathan):

1469.52

OpenStudy (studygurl14):

no

OpenStudy (leenathan):

that's not one of the answer choices

OpenStudy (studygurl14):

\(\Large\frac{1}{3}(3.14)(3)^2(6)\)

OpenStudy (leenathan):

i just need the answer

OpenStudy (sissyedgar):

C. 56.52 ft3.. ♥

OpenStudy (leenathan):

thanks

OpenStudy (sissyedgar):

Chu welcome.